Find the exact global maximum and minimum values of the function g(x)=7x−x2−15 if its domain is all real numbers.
1
Expert's answer
2020-03-03T14:18:07-0500

"f(x)=7x-x^2-15"

To find stationary points, the function is differentiated and equated to zero.

"f'(x)=7-2x"

"7-2x=0"

"7=2x, x=3.5"

There is one stationary point.

"7\times{3.5}-3.5^2-15=-2.75"

The critical point is (3.5,-2.75).

To determine whether it is a maximum or minimum, the second derivative is required.

"f''=-2"

Since the second derivative is negative, the stationary point is a maximum. Since there are no other maximum stationary points, (3.5,-2.75) is a global maximum. The function has no minimum point.